If your investment grows 2% each month for a whole year, what was the EFFECTIVE annual interest rate?
1
Expert's answer
2015-10-10T00:00:44-0400
If your investment grows 2% each month for a whole year, then the EFFECTIVE annual interest rate is Ie = (1 + i/n)^n - 1 = (1+0.02/12)^12 - 1 = 0.0202 = 2.02%